The adobe package^{∗}

Sebastian Rahtz S.Rahtz@elsevier.co.uk

1995/08/16

1 Introduction

This file contains suitable package files to load certain non-standard Adobe PostScript fonts. The font definition files and font metrics are available in both T1 and OT1 form in the companion CTAN fonts/psfonts collection.

2 The docstrip modules

1 ⟨∗garamond⟩

 2 \NeedsTeXFormat{LaTeX2e}
 3 \ProvidesPackage{garamond}[\filedate\space\fileversion\space
 4 Garamond PSNFSS2e package]
 5 \renewcommand{\rmdefault}{pgm}
 6 \renewcommand{\sfdefault}{pop}
 7 \renewcommand{\ttdefault}{pcr}

8 ⟨/garamond⟩ 9 ⟨∗basker⟩

10 \NeedsTeXFormat{LaTeX2e}
11 \ProvidesPackage{basker}[\filedate\space\fileversion\space
12 Baskerville PSNFSS2e package]
13 \renewcommand{\rmdefault}{pnb}

14 ⟨/basker⟩ 15 ⟨∗mtimes⟩

16 \NeedsTeXFormat{LaTeX2e}
17 \ProvidesPackage{mtimes}[\filedate\space\fileversion\space
18 Monotype Times PSNFSS2e package]
19 \renewcommand{\rmdefault}{mnt}

20 ⟨/mtimes⟩ 21 ⟨∗bembo⟩

22 \NeedsTeXFormat{LaTeX2e}
23 \ProvidesPackage{bembo}[\filedate\space\fileversion\space
24 Bembo PSNFSS2e package]

∗This file has version number 1.0, last revised 1995/08/16.

25 \renewcommand{\rmdefault}{pbb}
26 \renewcommand{\sfdefault}{pop}
27 \renewcommand{\ttdefault}{pcr}

28 ⟨/bembo⟩

Index

Numbers written in italic refer to the page where the corresponding entry is described, the ones underlined to the code line of the definition, the rest to the code lines where the entry is used.

F

\filedate . 3, 11, 17, 23

\fileversion 3, 11, 17, 23

N

\NeedsTeXFormat . . .

. . . . . 2, 10, 16, 22

P

                          \ProvidesPackage . .

. . . . . 3, 11, 17, 23

R

                          \renewcommand . . . . .

. 5–7, 13, 19, 25–27

                          \rmdefault 5, 13, 19, 25

S

                                                   \sfdefault . . . . . . 6, 26

\space …. 3, 11, 17, 23

T

                                                   \ttdefault . . . . . . 7, 27